OpenStudy (anonymous):

6 blue, 5 red and 4 green balls are arranged so that no 2 of the red balls are together. in how many ways can this be done if all balls of the same color are indistinguishable.

OpenStudy (anonymous):

answer key says 97020 O: i didn't get it though... i tried doing 15! over 5!6!4! - 11!over 6!4! O: O:

OpenStudy (anonymous):

I don't get that answer either. I got a lower number then that

OpenStudy (anonymous):

heeelp?? @ParthKohli ?@Hero @ash1122

OpenStudy (anonymous):

I'm no good at combinatrics since I'm still at it's early stages of learning, but do you have the correct answer?

OpenStudy (anonymous):

since each ball of the same colour is same first arrange 6 blue and 4 green balls which can be done in 10!/(6!*4!) now in between these there are 11 spaces which is to be filled with 5 red balls which can be done in (11*10*9*8*7)/5! ways =462 hence the required no of ways must be 10!/(6!*4!) *462=210*462=97020 (for calculation u may check once more )

OpenStudy (anonymous):

WOW THANK YOU!!!!!!! THANK YOUU!! <3

OpenStudy (anonymous):

welcome dear